Circular No. 7618

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

POSSIBLE SUPERNOVA IN NGC 3079 B. Swift, W. D. Li, and A. V. Filippenko, University of California at

Berkeley, report the discovery by LOTOSS (cf. IAUC 7514) of an apparent supernova in unfiltered images taken with the Katzman Automatic Imaging Telescope (KAIT): “The new object is located at α = 10^{h}01^{m}57^{s}.33, δ = +55^{o}41′14′′.6 (equinox 2000.0), which is 6′′.3 west and 25′′.4 north of the well-known starburst galaxy NGC 3079. Available photometry for the new object (measured after subtracting a template image observed on 1998 Nov. 21 UT): 2001 Apr. 17.2 UT, > 19.0; 25.2, 18.3; 26.2, 18.4; May 1.2, 18.1; 2.2, 17.7 (under bad conditions). Adopting a distance modulus of

30.75 for NGC 3079 from LEDA (http://cisr.univ-lyon1.fr/~leda/),

the absolute magnitude of the new star is only ∼ −12.7, similar to the type-IIn SN 1997bs (Van Dyk et al. 2000, PASP 112, 1532); this may be a superoutburst of a luminous blue variable rather than a genuine supernova. Other explanations for the apparent faintness of this new star is that it is intrinsically faint, or suffers from huge amount of extinction in the very dusty host galaxy, or both. A finding chart can be found at

http://astron.berkeley.edu/~bait/2001/ngc3079.html.”

SUPERNOVA 2001ad IN NGC 6373 R. Chornock, M. Modjaz, and A. V. Filippenko, University of Califor-

nia at Berkeley, report that inspection of a CCD spectrum (range 330–780 nm), obtained on Apr. 30 UT with the Shane 3-m reflector at Lick Observatory, reveals that SN 2001ad (cf. IAUC 7596, 7597) is actually an unusual ‘type-IIb’ supernova like SN 1987K (Filippenko 1988, A.J. 96, 1941) and SN 1993J (Filippenko et al. 1993, Ap.J. 415, L103): lines of He I with P-Cyg profiles are now prominent.

SUPERNOVA 2001ax IN ANONYMOUS GALAXY P. Nugent and L. Wang, Lawrence Berkeley National Laboratory, re-

port: “A spectrum (range 400-800 nm) obtained on Apr. 29.43 UT at the Kitt Peak 2.1-m telescope (+ GCAM) shows that SN 2001ax (cf. IAUC 7608) is a type-II supernova more than a month after peak brightness. The redshift of the supernova is ∼ 6000 km/s. The supernova shows a broad Hα emission feature, a strong Ca II infrared-triplet P-Cyg profile, and Fe II absorption lines.”

2001 May 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green